# MAFE: Multi-Agent Fair Environments for Decision-Making Systems

### v1

@@ Abstract
Fairness constraints applied to machine learning (ML) models in static contexts have been shown to potentially produce adverse outcomes among demographic groups over time. To address this issue, emerging research focuses on creating fair solutions that persist over time. While many approaches treat this as a single-agent decision-making problem, real-world systems often consist of multiple interacting entities that influence outcomes. Explicitly modeling these entities as agents enables more flexible analysis of their interventions and the effects they have on a system’s underlying dynamics. A significant challenge in conducting research on multi-agent systems is the lack of realistic environments that leverage the limited real-world data available for analysis. To address this gap, we introduce the concept of a Multi-Agent Fair Environment (MAFE) and present and analyze three MAFEs that model distinct social systems. Experimental results demonstrate the utility of our MAFEs as testbeds for developing multi-agent fair algorithms.
Machine Learning, ICML
@@ 1 Introduction
As artificial intelligence (AI) and machine learning (ML) become increasingly embedded in everyday life, concerns about their potential to exacerbate social biases have grown (Sweeney, 2013; Angwin et al., 2016; Larson et al., 2016; Buolamwini & Gebru, 2018). Algorithms aimed at addressing these issues, known as algorithmic fairness, must minimize bias, integrate seamlessly into decision-making systems, and provide meaningful interventions against systemic disparities. Early approaches focused on mitigating static biases, such as group (Kamiran & Calders, 2012; Hardt et al., 2016), individual (Dwork et al., 2012), and causal (Kusner et al., 2017; Coston et al., 2020) bias.
However, studies have shown that these methods can fail to address, or even worsen, the differential harms that emerge as decisions unfold over time (Liu et al., 2018; D’Amour et al., 2020). For example, in healthcare, a treatment algorithm may allocate resources equally at diagnosis, ensuring the fairness of an isolated decision. However, as time progresses, differing demographic factors—such as access to follow-up care or socioeconomic disparities—can lead to worsened outcomes for certain groups, highlighting the need for interventions that can continuously address and mitigate these disparities dynamically over time.
To address this issue, recent works have reframed fairness as a sequential decision-making problem, often modeled using Markov Decision Processes (MDPs) (Yin et al., 2024; Xu et al., 2024) or structural causal models (Hu & Zhang, 2022). While these methods often model the decision-making process of an agent at a single decision point within the system, many real-world systems involve multiple decision points that cumulatively impact the entire system. For example, in healthcare, one decision point might be providing insurance to individuals, while another consists of ensuring they have access to timely and quality healthcare services, such as a hospital bed, when such resources are limited. If some individuals receive insurance but live in areas with restricted healthcare access, disparities in outcomes may still emerge. Comprehensive strategies—such as insurance assistance or access to healthcare—are thus essential to address systemic disparities and ensure fairness across decision stages.
(a) MAFE Diagram.
(b) Healthcare MAFE Example.
Figure 1: Illustration of our MAFE definition. (a) A diagram capturing the elements of our MAFE. The actions produce by the model(s) are imported to the environment to be taken by environmental agents. This leads to state transition within the environment that produces a set collection of observations, rewards, and fairness components for each agent which are output by the environment for the model(s) to use to produce actions in the next time step. (b) An example illustrating this process for a healthcare MAFE that particularly captures how the component functions can be used to construct measures of rewards and fairness.
This underscores the need for realistic frameworks to study fairness dynamics in multi-agent systems over time. However, the lack of realistic multi-agent environments for studying such systems presents a major obstacle for developing and evaluating effective algorithms for addressing disparities in various social contexts. To bridge this gap, we propose a framework for Multi-agent Fair Environments (MAFEs) that simulate modular social systems.
Summary of Contributions: We introduce MAFEs as a novel framework for analyzing fairness in multi-agent systems and present three benchmark MAFEs designed to simulate critical social domains: loan processing, healthcare, and higher education. To demonstrate the versatility of MAFEs and facilitate empirical evaluation, we present a cooperative use case. Within this use case, we define success metrics tailored to algorithms operating in these MAFEs and showcase the adaptability of a multi-agent reinforcement learning (MARL) algorithm in this setting. Additionally, we conduct thorough experimental analysis of our MAFEs and offer insights into their dynamics.
@@ 2 Related Works
To overcome the limitations of static fairness formulations, several approaches have re-framed fairness as a dynamic systems problem. Effort-based fairness analyzes the differing efforts required by groups to achieve outcomes (Heidari et al., 2019; Guldogan et al., 2022), while causal models use structural causal models and interventions to introduce fairness (Hu et al., 2020; Hu & Zhang, 2022). Another approach incorporates fairness within dynamic systems through reinforcement learning (RL), with early work using multi-armed bandits (Joseph et al., 2016) and recent efforts employing Markov Decision Processes (MDPs). Puranik et al. (Puranik et al., 2022) introduce the Fair-Greedy policy in an admissions case study, balancing applicants’ scores with group proportions. Yin et al. (Yin et al., 2024) frame the long-term fairness RL problem to maximize profits while minimizing unfairness, measured by regret and distortion. To address temporal bias, Xu et al. (Xu et al., 2024) propose a fairness measure based on the ratio-after-aggregation and modify the proximal policy optimization algorithm (PPO) to satisfy this constraint. Though these works reduce temporal disparities, they do not analyze their source. Deng et al. (Deng et al., 2024) use causal analysis to trace sources of inequality over time. While these works extend static fairness to long-term outcomes, Hu et al. (Hu et al., 2023) argue that long-term fairness should focus on the convergence of input feature distributions, proposing a PPO variant with pre-processing and regularization to balance short- and long-term fairness.
@@
In systems with multiple decision-making entities, modeling fairness explicitly across agents becomes crucial for understanding their interventions and their effects on system dynamics. Several studies have explored fairness in multi-agent contexts. Jiang and Lu (Jiang & Lu, 2019) introduce the Fair-Efficient Network, a hierarchical RL model where homogeneous agents aim to balance fairness and efficiency. Zheng et al. (Zheng et al., 2022) use two-level deep RL to design agents that reduce income inequality via taxation and redistribution, with equity measured by the Gini Index. Reuel and Ma (Reuel & Ma, 2024) provide a survey on fairness in RL, covering both single- and multi-agent systems. They highlight key gaps, such as fairness in RL from human feedback, and emphasize the challenges of ensuring fairness in dynamic real-world environments, which underscores the need for realistic simulation environments.

### v2

# MAFE: Enabling Equitable Algorithm Design in Multi-Agent Multi-Stage Decision-Making Systems

@@ Abstract
Algorithmic fairness is often studied in static or single-agent settings, yet many real-world decision-making systems involve multiple interacting entities whose multi-stage actions jointly influence long-term outcomes. Existing fairness methods applied at isolated decision points frequently fail to mitigate disparities that accumulate over time. Although recent work has modeled fairness as a sequential decision-making problem, it typically assumes centralized agents or simplified dynamics, limiting its applicability to complex social systems. We introduce MAFE, a suite of Multi-Agent Fair Environments designed to simulate realistic, modular, and dynamic systems in which fairness emerges from the interplay of multiple agents. We demonstrate MAFEs in three domains—loan processing, healthcare, and higher education—supporting heterogeneous agents, configurable interventions, and fairness metrics. The environments are open-source and compatible with standard multi-agent reinforcement learning (MARL) libraries, enabling reproducible evaluation of fairness-aware policies. Through extensive experiments on cooperative use cases, we demonstrate how MAFE facilitates the design of equitable multi-agent algorithms and reveals critical trade-offs between fairness, performance, and coordination. MAFE provides a foundation for systematic progress in dynamic, multi-agent fairness research.
@@ 1 Introduction
As machine learning (ML) systems increasingly shape decisions in critical domains, such as lending, healthcare, and education, concerns have intensified about their potential to exacerbate social inequities (Sweeney, 2013; Angwin et al., 2016; Larson et al., 2016; Buolamwini and Gebru, 2018). The field of algorithmic fairness seeks to design interventions that not only mitigate bias at the point of decision but also prevent disparities from compounding over time.
While early approaches focused on static definitions of fairness—targeting group-level (Kamiran and Calders, 2012; Hardt et al., 2016), individual-level (Dwork et al., 2012), and causal (Kusner et al., 2017; Coston et al., 2020) biases—these criteria often fall short in dynamic settings. For example, a healthcare system that ensures equal treatment at diagnosis may still produce inequitable long-term outcomes if certain populations face barriers to follow-up care (Liu et al., 2018; D’Amour et al., 2020). Addressing such evolving disparities demands frameworks that capture sequential decisions and their cumulative effects.
Recent works model fairness through sequential lenses, using Markov Decision Processes (MDPs) (Yin et al., 2024; Xu et al., 2024) or structural causal models (Hu and Zhang, 2022). However, these methods generally assume a single decision-maker operating in isolation. In contrast, real-world systems are multi-actor: insurers, hospitals, and government agencies jointly influence population health; schools, employers, and regulators together shape educational equity. Capturing such systems requires a shift to multi-agent formulations where fairness is not a property of one decision, but of distributed interactions across agents.
Yet progress in this direction is constrained by the absence of realistic, modular environments for evaluating fairness in multi-agent systems. Existing platforms typically assume centralized control, lack support for heterogeneity across agents, or oversimplify social dynamics, limiting their utility for fairness-aware algorithm design.
To bridge this gap, we introduce MAFE  [^1] We release our anonymized codebase at https://anonymous.4open.science/r/MAFE_Environments-88CA/README.md—a benchmark suite of Multi-Agent Fair Environments for modeling dynamic decision-making systems where fairness arises from the interactions of multiple agents. Each MAFE is a modular and extensible simulation of a social system, featuring heterogeneous agents, configurable disparities, and evolving population dynamics. Designed to support algorithm development and empirical evaluation, MAFE offers a principled testbed for fairness-aware multi-agent learning.
Summary of Contributions.
By introducing MAFE as a benchmark suite of Multi-Agent Fair Environments for evaluating fairness-aware policies in dynamic, multi-agent settings, we provide the following key contributions:
  • Framework and Benchmarks. We propose the MAFE framework and instantiate three open-source environments—MAFE-Loan, MAFE-Health, and MAFE-Edu—that model equity challenges across social domains.
  • Fairness-Aware Modeling and Evaluation. We define a cooperative use case and formalize fairness-aware optimization objectives that capture long-term system equity. We introduce diagnostic metrics and illustrate how MARL algorithms can be adapted to equity-driven reward structures through a representative implementation.
  • Empirical Validation. We evaluate the behavior of a representative MARL algorithm in MAFE environments, offering reproducible baselines and highlighting trade-offs between fairness and utility.
@@ 2 Related Works
To overcome the limitations of static fairness formulations, several approaches have re-framed fairness as a dynamic systems problem. Effort-based fairness analyzes the differing efforts required by groups to achieve outcomes (Heidari et al., 2019; Guldogan et al., 2022), while causal models use structural causal models and interventions to introduce fairness (Hu et al., 2020; Hu and Zhang, 2022). Another approach incorporates fairness within dynamic systems through reinforcement learning (RL), with early work using multi-armed bandits (Joseph et al., 2016) and recent efforts employing Markov Decision Processes (MDPs). (Puranik et al., 2024) introduce the Fair-Greedy policy in an admissions case study, balancing applicants’ scores with group proportions. (Yin et al., 2024) frame the long-term fairness RL problem to maximize profits while minimizing unfairness, measured by regret and distortion. To address temporal bias, (Xu et al., 2024) propose a fairness measure based on the ratio-after-aggregation and modify the proximal policy optimization algorithm (PPO) to satisfy this constraint. Though these works reduce temporal disparities, they do not analyze their source. (Deng et al., 2024) use causal analysis to trace sources of inequality over time. While these works extend static fairness to long-term outcomes, (Hu et al., 2023) argue that long-term fairness should focus on the convergence of input feature distributions, proposing a PPO variant with pre-processing and regularization to balance short- and long-term fairness.
@@
In systems with multiple decision-making entities, modeling fairness explicitly across agents becomes crucial for understanding their interventions and their effects on system dynamics. Several studies have explored fairness in multi-agent contexts. (Jiang and Lu, 2019) introduce the Fair-Efficient Network, a hierarchical RL model where homogeneous agents aim to balance fairness and efficiency. (Zheng et al., 2022) use two-level deep RL to design agents that reduce income inequality via taxation and redistribution, with equity measured by the Gini Index. (Reuel and Ma, 2024) provide a survey on fairness in RL, covering both single- and multi-agent systems. They highlight key gaps, such as fairness in RL from human feedback, and emphasize the challenges of ensuring fairness in dynamic real-world environments, which underscores the need for realistic simulation environments.
Figure 1: MAFE Diagram. Model(s) produce actions that are imported to the environment and taken by agents. This leads to state transition within the environment that produces a set collection of observations, rewards, and fairness components for each agent which are output by the environment for the model(s) to use to produce actions in the next time step.